OpenStudy (anonymous):

1 ---% × 10 = _____ 5 Select one of the options below as your answer:A. .02 B. .05 C. .15 D. 2

myininaya (myininaya):

\[\frac{1}{5} \cdot \frac{1}{100} \cdot 10=\frac{1}{5} \cdot \frac{10}{100}=\frac{1}{5} \cdot \frac{1}{10}=\frac{1}{50}\]

myininaya (myininaya):

You guys don't just put the answers.

myininaya (myininaya):